# Dynamical stability for dense patterns in discrete attractor neural networks

### v1

@@ Introduction.
with a new cross-correlation term $c_{rf0}^{\mu}=\left\langle\delta r^{\mu}\delta f_{0}\left(r^{\mu}\right)\right\rangle$. Empirically, it is straightforward to identify the eigenvector with the highest cross-correlation to the memory pattern (associated eigenvalue marked in orange at Fig. 2a), systematically assess the validity of the estimator $\lambda_{\mathrm{mem}}$ (Fig. 2e) and find cases where they determine the dynamical stability of the Jacobian (Fig. 2b, right column).
We note an interesting symmetry: the bulk-related instability $\lambda_{\mathrm{bulk}}\geq 0$ when $c_{rf}\geq c_{rr}$, and the pattern-specific instability $\lambda_{\mathrm{mem}}^{\mu}\geq 0$ when $c_{rf0}^{\mu}\geq c_{rr}$. The former was derived by neglecting pattern-specific statistical dependency between Jacobian rows, while the latter is a direct result of this dependency.
Importantly, the two outliers $\lambda_{\mathrm{ave}},\lambda_{\mathrm{mem}}$ depend only on the pattern statistics and the activation function, but not on $\alpha$, i.e., on the number of stored patterns, so our estimate for the critical load for stability:
$\alpha_{\mathrm{S}}=\alpha_{\mathrm{S}}^{\mathrm{bulk}}\Theta\left(-\lambda_{\mathrm{mem}}\right)\Theta\left(-\lambda_{\mathrm{ave}}\right)\Theta\left(\tau_{\mathrm{mem}}-0.9\right)$ (17)
To show the qualitative and quantitative predictions of the theory more systematically, we numerically analyzed the critical load for stability for different combinations of pattern statistics CV, activation function smoothness $\sigma$, and exponent $n$. Fig. 3 present the results as phase diagrams in the $\sigma$ vs $n$ plane. Critical stability is zero in many areas on this plane, with the stable regime defined by the intersection of 3 conditions: (1) $\lambda_{\mathrm{bulk}}>0$ prevents stability at most of the sublinear regime; (2) $\lambda_{\mathrm{mem}}>0$ prevents stability at most of the supralinear regime; (3) $\tau_{\mathrm{mem}}<0.9$ at large values of $\sigma$, invalidating the assumption of Eq. 15; so instability is not captured by Eq. 16. We thus predict a surprising mechanism for instability for $n>1$, due to $\lambda_{\mathrm{mem}}>0$ (Fig. 3). Intuitively, in this regime, a perturbation in the direction of the memory pattern, or scaling it up by a factor, will be amplified due to a supralinear exponent and diverge.
Figure 3: Phase diagram of stability. (a-e) The predicted value of $\alpha_{\mathrm{S}}$ (color coded) for different combinations of I/O exponent (x-axis) and I/O smoothness (y-axis, log-scale). A plus sign marks the location of the maxima. Pattern variation level (CV) is indicated at each panel. Boundaries of the different regimes are marked: $\lambda_{\mathrm{bulk}}>0$ (blue line), $\lambda_{\mathrm{mem}}>0$ (orange line), $\tau_{\mathrm{mem}}<0.9$ (red line). (a) also overlays a grid of numerical results: yellow disk marks stability (with size-coded value), gray cross marks stability below detectable range (most crosses removed to avoid clutter).
Figure 4: Optimal parameters. (a-b) The optimal exponent (a) or smoothness (b), at different pattern variation levels (x-axis). (c) The critical load for stability using optimal parameters, at different pattern variation levels (x-axis). (d) Measures of the weights asymmetry (dark) and dynamics non-normality (light) at different pattern variation levels (x-axis). The asymmetry index is $\left\|\mathbf{W}_{\mathrm{asym}}\right\|_{\mathrm{F}}/\left(\left\|\mathbf{W}_{\mathrm{sym}}\right\|_{\mathrm{F}}+\left\|\mathbf{W}_{\mathrm{asym}}\right\|_{\mathrm{F}}\right)$ for the symmetric and anti-symmetric parts of the weights $\mathbf{W}$. The non-normality index is Henrici’s deviation from normality index [22, 23] for the Jacobian $\mathbf{J}$, $\sqrt{\left\|\mathbf{J}\right\|_{\mathrm{F}}^{2}-\sum_{i}\left|\lambda_{i}\right|^{2}}/\left\|\mathbf{J}\right\|_{\mathrm{F}}$. Presenting the median value across choices of exponent and smoothness.
Optimal parameter.
The overall comparison of theory and numerical experiments is presented in Fig. 8. While the match is not perfect, Fig. 3 explains the sources of the discrepancy. The points with numerical results close to zero but substantial prediction from theory, or vice versa, all come from the $n\approx 1$ region, where any imprecision in estimating the boundary $\lambda_{\mathrm{mem}}=0$ changes the prediction between an order-one value and 0.
Fig. 4a-b shows the optimal exponent and smoothness for different levels of pattern variation. Our analysis predicts substantial stability only in the near-linear region $n\approx 1$, and at a finite value of $\sigma$. The benefits of threshold-linear activation are also observable for sparse patterns (Fig. 6e), where it is optimal for $\textrm{CV}\geq 1$. This is consistent with previous literature on storage of graded memory patterns using the ‘Hebbian’ approach, which focused on this activation function [5, 12] and with the machine learning literature using similar activations [24], but contrasts with other approaches which advocates for the benefits of supralinear activation [25, 26].
@@
Theory further predicts improvement of stability with pattern variation (Fig. 4c). Those high-variance memory patterns are sparse-like, with many near-zero values and few large activations, so the improved capacity we report for those echoes the benefit of sparseness in our results (Fig. 6e) and previous works [5, 4]. Such sparse-like patterns are also consistent with the known phenomenology of neurons in the brain [27, 28].
To test if the optimal weights used here may coincide with those assumed by the ‘Hebbian’ approach, we measured the weight matrix asymmetry and the deviation from normality of the dynamics around the fixed points (Fig. 4d). Both measures tend to increase with variation levels, indicating that a high critical load for stability is likely achieved for asymmetric weights, and leads to dynamics distinct from those of energy-based models.
@@
The resulting optimal threshold is usually negative in our theory for dense patterns (Eq. 10), and also empirically for sparse patterns (Fig. 6c). A negative threshold is uncommon in the literature (a notable exception is the Willshaw model [29]) and contrasts with the prevailing view in the field [17], which holds that the construction of multiple distinct and stable activity patterns requires global inhibition and selective recurrent excitation.
@@ Discussion.
Our analysis highlights the conditions for memory pattern stability in the auto-associative memory task. The resulting theory is rich and makes many testable predictions, most notably a stability phase transition at a number of patterns which is proportional to the number of neurons, and is below the capacity phase transition, which does not account for stability. It constrains the range of usable activation thresholds and predicts the optimal single-neuron activation function for different memory pattern statistics. Most notably, we establish that near-linear activation and finite noise levels are optimal for stable recall of dense memory patterns.
@@
Those insights are relevant to neuroscience and cannot be derived from the non-biological, energy-based, auto-associative memory networks commonly used in the field. This letter opens the door for designing networks with biological-relevant dynamics capable of storing many patterns as stable fixed points, with simple optimization. Future research may establish if biological-like local learning can be used for such optimization.
@@
Many areas of science study dynamical systems where fixed points and their dynamic stability may be of interest [30]. Theoretically, this question may be analyzed through the eigenvalue spectrum of the dynamics’ Jacobian at the fixed point. Full characterization of the spectrum is possible for many classes of random matrices [31, 32, 33, 34, 35, 36, 37] and exhibits a universality property [38]. Many random dynamical systems exhibit a stability phase transition when the largest real value in the eigenvalues spectrum crosses zero [30, 39, 36, 40, 41]. Those random matrices can have an exponential number of fixed points, with stability which depends on a global criterion [42, 35, 43]. In this letter, we go beyond random matrices (and low-rank perturbations thereof [44, 45]) to describe fixed-point stability in a system resulting from a learning or an optimization process. In such systems, the connection weight matrix (and subsequently the Jacobian) is defined in terms of the constraints it satisfies or the objective function it minimizes. The stability of fixed points in a learned network could be previously analyzed only for very simple cases, such as a scalar output [46]. Even for the heavily studied Hopfield model, the stability of patterns remains unclear, and memory patterns are not, in general, fixed points of the dynamics [2]. Our analysis of the eigenvalue spectrum of the Jacobian may be applicable to non-random systems in other fields, when the optimal connectivity is given by a pseudo-inverse rule. Admittedly, this analysis needs to be complemented with a problem-specific approach to characterize outliers.

### v3

@@ Introduction.
where $c_{\mathrm{r\phi}}=\left\langle\delta r\,\delta\phi\!\left(r\right)\right\rangle$, $c_{\mathrm{\phi\phi}}=\left\langle\delta\phi\!\left(r\right)^{2}\right\rangle$ are additional (co)variance terms, and the assumption on which the analysis of $\lambda_{\mathrm{mem}}$ is based can be expressed as $\tau_{\mathrm{mem}}\approx 1$. To empirically test the relevance of $\lambda_{\mathrm{mem}}^{\mu}$ (Eq. 15) for determining the stability of $\mathbf{J}_{\mu}^{*}$, it is straightforward to identify the eigenvector of $\mathbf{J}_{\mu}^{*}$ associated with the highest $\tau_{\mathrm{mem}}^{\mu}$, and compare the corresponding eigenvalue to $\lambda_{\mathrm{mem}}$ (Fig. 2a,e). Indeed, we find cases where $\lambda_{\mathrm{mem}}^{\mu}$ determines the dynamical stability of $\mathbf{J}_{\mu}^{*}$ (Fig. 2b, right column).
For characterizing the transition from the stable to the unstable regime, we analyze the case when exactly half of the patterns stored in a network are stable. This means that we focus on the median $\lambda_{\mathrm{ave}}^{\mu}$ and $\lambda_{\mathrm{mem}}^{\mu}$ across patterns. As the expected value of this median is the expectation of an individual such eigenvalue, stability is determined by the zero crossings of the expectations we computed above (Eqs. 12 and 15). As neither of these expectations depend on the number of stored patterns, we can estimate the critical load for stability as:
$\alpha_{\mathrm{S}}=\alpha_{\mathrm{S}}^{\mathrm{bulk}}\,\Theta\!\left(-\lambda_{\mathrm{ave}}\right)\,\left[1-\Theta\!\left(\lambda_{\mathrm{mem}}\right)\,\Theta\!\left(\tau_{\mathrm{mem}}-0.95\right)\right]$ (16)
where $\tau_{\mathrm{mem}}>0.95$ is an arbitrary choice for the domain where we consider the prediction about $\lambda_{\mathrm{mem}}$ to apply.
We first focus on the zero crossing of $\alpha_{\mathrm{S}}$ as a function of the parameters, i.e. the parameter regime where stability is possible at all. For this, we analyze the zero crossing of each of the three terms of Eq. 16 separately by solving the corresponding inequalities: $\alpha_{\mathrm{S}}^{\mathrm{bulk}}>0$, $\lambda_{\mathrm{ave}}<0$, and $\lambda_{\mathrm{mem}}<0$ (assuming $\tau_{\mathrm{mem}}>0.95$). The dependence of each of these zero crossings on $\theta$ is relatively straightforward. The zero crossing of $\alpha_{\mathrm{S}}^{\mathrm{bulk}}$ does not depend on $\theta$ (although the exact value of $\alpha_{\mathrm{S}}^{\mathrm{bulk}}$ when it is greater than zero does, Details). Both $\lambda_{\mathrm{ave}}$ and $\lambda_{\mathrm{mem}}$ depend on $\theta$ linearly (Eq. 12, and Details), and so the condition for their zero crossing can be expressed respectively as
$\theta_{0}^{\mathrm{ave}}$ $=\frac{\langle r\rangle}{\left\langle d\!\left(r\right)\right\rangle}-\left\langle g^{-1}\!\left(r\right)\right\rangle$ (17)
$\text{and }\theta_{0}^{\mathrm{mem}}$ $=\frac{\left\langle\delta r\,\delta r\right\rangle-\left\langle\delta r\,\delta d\!\left(r\right)\,g^{-1}\!\left(r\right)\right\rangle}{\left\langle\delta r\,\delta d\!\left(r\right)\right\rangle}$ (18)
where $d\!\left(r\right)=g^{\prime}\!\left(g^{-1}\!\left(r\right)\right)$, and the RHSs of Eqs. 17 and 18 implicitly depend on the other parameters (Details).
Phase diagrams.
In order to gain insights into the complex dependence of the zero crossings of the three terms of Eq. 16 on the combination of parameters (pattern statistics CV, activation function smoothness $\sigma$, exponent $n$, and threshold $\theta$), we solved the corresponding equations systematically over a range of parameter combinations, and compared these predictions with numerically obtained values of $\alpha_{\mathrm{S}}$. Fig. 3a presents the resulting phase diagrams of $\alpha_{\mathrm{S}}$ for different pairs of parameters when the third parameter is kept fixed, showing where $\alpha_{\mathrm{S}}^{\mathrm{bulk}}$, $\lambda_{\mathrm{ave}}$, and $\lambda_{\mathrm{mem}}$ (the three main determinants of $\alpha_{\mathrm{S}}$ according to Eq. 16) are predicted to cross zero (blue, green, and orange lines, respectively) as well as the numerically computed values of $\alpha_{\mathrm{S}}$ (grayscale map). We note that $\lambda_{\mathrm{ave}}$ constrains stability only for $\theta>1,n\approx 1$ (Fig. 3a middle panel), while the upper bound on the threshold $\theta$ is due to the zero crossing of $\lambda_{\mathrm{mem}}$ (Eq. 18). The superior stability near $n=1$ is explained by noting that while $\lambda_{\mathrm{bulk}}$ prevents stability at most of the sublinear regime, where $c_{\mathrm{rf}}\geq c_{\mathrm{rr}}$, $\lambda_{\mathrm{mem}}$ prevents stability at most of the supralinear regime, where $c_{\mathrm{r\phi}}\geq c_{\mathrm{rr}}$. Interestingly, the former is derived by neglecting the pattern-specific statistical dependency between the rows of the Jacobian, while the latter is a result of this dependency.
Furthermore, we can now better understand the limitations of the theory: the zero crossing of $\lambda_{\mathrm{mem}}$ yields a discontinuous fall of $\alpha_{\mathrm{S}}$ from a finite value to 0 (Eq. 16), and in many cases the maxima is achieved at this boundary. Thus, deviations of the empirical value of this outlier would cause large prediction errors in that boundary. Furthermore, in two regimes – large $\sigma$ and low $\theta$ – we observe regions where the assumption on the memory-pattern related outlier is violated, and the empirical value of the outlier is larger than theory predicts, so the empirical $\alpha_{\mathrm{S}}$ might be zero while theory predicts otherwise.
To complement these results, Fig. 3b presents similar phase diagrams, but where the third parameter is optimized for each pair of systematically varied parameters, rather than kept fixed. The results remain qualitatively similar to those obtained above (Fig. 3a). This view highlights the superiority of $n=1$ and $\sigma=1$ in terms of the resulting number of stable memories, which is further demonstrated across all CV values in Fig. 3c-d for $n$ and $\sigma$, respectively. On the other hand, the effect of $\theta$ seems minor, as long as it is below an upper bound.
Figure 3: Phase diagrams for the critical load for stability. (a-b) Numerical simulations results for the critical load for stability $\alpha_{\mathrm{S}}$ (grayscale map) using different choices of I/O exponent $n$, smoothness $\sigma$, and threshold $\theta$, and theoretical predictions for the zero-crossings of the different terms in Eq. 16, i.e. $\alpha_{\mathrm{S}}^{\mathrm{bulk}}$, $\lambda_{\mathrm{ave}}$, and $\lambda_{\mathrm{mem}}$ (blue, green, and orange curves, respectively). $\lambda_{\mathrm{mem}}$ is considered only in the domain where $\tau_{\mathrm{mem}}>0.95$. In (a), two parameters are varied (axes), while the third parameter (title) is kept fixed. In (b), the third parameter is optimized for each combination of the two parameters that are being varied. $\mathrm{CV}=4$ in all cases shown. (c) The maximal value of $\alpha_{\mathrm{S}}$ for each value of $n$ (x-axis) and $\mathrm{CV}$ (color coded). (Parameters $\sigma$ and $\theta$ are optimized for each combination of $n$ and $\mathrm{CV}$.) (d) The maximal value of $\alpha_{\mathrm{S}}$ for each value of $\sigma$ (x-axis) and $\mathrm{CV}$ (color coded as in c). (Parameters $n$ and $\theta$ are optimized for each combination of $n$ and $\mathrm{CV}$.)
Optimal parameters.
@@
Fig. S4 presents the maximal achievable value for the stability critical load as a function of pattern statistics, as well as the optimal parameters where the optimum is achieved for various choices of $\left<r\right$>. Our theory predicts improvement of stability with pattern variation (Fig. 3c-d, Fig. S4a). Those high-variance memory patterns are sparse-like, with many near-zero values and few large activations, so the improved capacity we report for those echoes the benefit of sparseness in our results (Fig. S2a,e) and previous works [5, 4]. Such sparse-like patterns are also consistent with the known phenomenology of neurons in the brain [23, 24]. As noted above, such patterns require highly asymmetric weights and lead to dynamics distinct from those of energy-based models (Fig. 1i).
Our analysis predicts substantial stability only in the near-linear region $n\approx 1$ (Fig. 3c, Fig. S4b). The benefits of threshold-linear activation are also observable for sparse patterns (Fig. S2e), where it is optimal for $\textrm{CV}\geq 1$. This is consistent with previous literature on storage of graded memory patterns using the ‘Hebbian’ approach, which focused on this activation function [5, 12] and with the machine learning literature using similar activations [25], but contrasts with other approaches that advocate for the benefits of supralinear activation [26, 27].
Numerical results demonstrate that the optimal choice of $\sigma$ is a finite value (Fig. 3d, Fig. S4c). Theory does not predict this value well due to the abovementioned imprecisions in the estimation of the zero crossing of $\lambda_{\mathrm{mem}}$.
@@
In our analysis, the optimal threshold is negative. This is supported by the theory bounding the threshold from above (Fig. 3a-b), and is evident empirically at the optimal parameters for dense patterns (Fig. S4d), as well as for sparse patterns (Fig. S2c). A negative threshold implies that neurons are biased toward activity and are held back by recurrent connections which are dominated by inhibition (i.e. $m<0$), as shown for the near-optimal values $n=1,\sigma=1$ at Fig. 1g. This contrasts with previous proposals for how recurrent neural networks can perform associative memory (a notable exception is the Willshaw network [28]) and with the prevailing view in the field [17], according to which the construction of multiple distinct and stable activity patterns requires global inhibition and selective recurrent excitation.
@@ Discussion.
Our analysis highlights the conditions for memory pattern stability in attractor networks performing auto-associative memory recall. The resulting theory makes several testable predictions, most notably a stability phase transition at a number of patterns which is proportional to the number of neurons, below the capacity phase transition, which does not account for stability. It constrains the range of possible choices for single-neuron activation functions and establishes that near-linear activation, a negative threshold, and non-zero noise levels are optimal for the stable recall of dense memory patterns.
@@
The predictions above are relevant to neuroscience and cannot be derived from the non-biological, energy-based auto-associative memory networks commonly used in the field. This letter paves the way for designing networks with biologically relevant dynamics capable of storing a large number of patterns as stable fixed points by standard optimization approaches. In particular, optimization only needs to consider the existence of fixed points, not their stability, as the latter is guaranteed under the stability phase transition. Future research will need to establish if biologically plausible local learning rules, rather than optimization, can achieve similar results.
@@
Several areas of science study dynamical systems where fixed points and their dynamic stability may be of interest [29]. This question is analyzed theoretically, as in our case, through the eigenvalue spectrum of the dynamics’ Jacobian at fixed points. Full characterization of the spectrum is possible for a variety of random matrix ensembles [30, 31, 32, 33, 34, 35, 36] and exhibits a universality property [37]. Many random dynamical systems (whose local linearization induces a random matrix ensemble) exhibit a stability phase transition as some parameter of the ensemble is increased [29, 38, 35, 39, 40]. Such random dynamical systems can have an exponential number of fixed points, with stability that depends on a global criterion [41, 34, 42]. In this letter, we were able to go beyond random matrices (or low-rank perturbations thereof [43, 44]) to describe fixed-point stability in a system resulting from an optimization process, where the connection weight matrix (and subsequently the Jacobian) has considerable structure defined by the objective function it minimizes. The stability of fixed points in an optimized network was previously described analytically only for simple cases, when storing a single pattern of activity by optimizing the weights of a single feedback neuron (i.e. a rank-1 perturbation of the weights) [45], or a very small number of patterns using a rank-2 perturbation [46]. Our work provides conditions for the dynamical stability of networks optimized to store a large number of memory patterns.
Our analysis of the eigenvalue spectrum of the Jacobian may be applicable to many other systems where an optimal solution is given by a pseudo-inverse, as in multivariate linear regression problems $S^{*}=\arg\min_{S}\|A-BS\|_{F}$ for rectangular matrices $A$, $B$ and a square matrix $S$. Those arise naturally in many areas of science, such as in the case of time series analysis where $A$ and $B$ are time-lagged “Trajectory Matrices”, useful for the analysis of non-linear dynamics (in “Singular Spectrum Analysis” [47, 48]), for Multivariate Calibration in Chemometrics (in “Principal Component Regression” [49]) and in the study of fluid dynamics (in “Dynamic Mode Decomposition” [50]). Our approach, in particular deriving expressions similar to Eq. 22 with appropriate assumptions, should allow the analysis of the optimal solution’s spectrum and its stability in this broad range of problems.